OpenStudy (hihi67):

i have some algebra equations that i have done. can you check them? (i will post them soon).

OpenStudy (hihi67):

a) k/3 + k/2 =

OpenStudy (hihi67):

b) 4/5 x - 1/10 x

OpenStudy (hihi67):

c) 3/8 z - 1/4 z + 1/2 z

OpenStudy (hihi67):

d) s/5 - s/3 + s

OpenStudy (hihi67):

e) 1 1/2 m - 3/4m + m/8

OpenStudy (hihi67):

and.....

OpenStudy (hihi67):

f) -4a/5 + 1/10a - 3a

OpenStudy (hihi67):

tell me what answers you get, and then i will see if i have the same. pleasee....?

OpenStudy (agreene):

a) \[\frac{5k}6\] b) \[\frac7{10x}\] c) \[\frac5{8z}\] d) \[\frac{13s}{15}\] e) \[\frac{m}{8}-\frac{1}{4m}\] f) \[\frac1{10a}-\frac{19a}5\]

OpenStudy (hihi67):

okay... well for a b and c i have the same. for d i have - 17/15s. for e i have 6/8m. for f i have - 3 and 9/10a

OpenStudy (hihi67):

this is how i figured it out. d:

OpenStudy (agreene):

from d down it was a bit hard to know what the question was--since you didn't use parenthesis.

OpenStudy (hihi67):

no () in book either.

OpenStudy (hihi67):

ust a min.

OpenStudy (agreene):

d could have been: \[\frac{s}5 - \frac{s}3 + s\] which is what i assumed, or \[\frac{s}5-\frac{s}{3+s}\] which would be a harder problem, lol

OpenStudy (hihi67):

\[s/5 - s/3 + s : \space \space \space \space \space \space \space \space \space s-5-s/3 = 3s/15 - 5s/15 \space \space \space that equals -2s/15 \space \space \space \space \space -2s/15 + s = \]

OpenStudy (hihi67):

thats d

OpenStudy (hihi67):

dang cant see it.

OpenStudy (hihi67):

never mind...

OpenStudy (hihi67):

bye thanks

OpenStudy (agreene):

lol dont give up so quickly :P

OpenStudy (hihi67):

ooooh I KNOW!

OpenStudy (hihi67):

stupid "order of operations"

OpenStudy (hihi67):

yep, thanks! you really actually did help a lot!!!

OpenStudy (agreene):

for d: I find it easier to start by factoring out the variables (makes it easier for me to see): s/5 - s/3 +s \[\frac15(s) - \frac13(s)+ s\] combine your fractions: \[-\frac2{15}(s)+s\] combine those fractions (convert the +s to +15s/15 so its the same denominator) \[\frac{15}{15}(s) - \frac2{15}(s) = \frac{15-2}{15}(s)==\frac{13}{15}(s)\] or \[\frac{13s}{15}\] thats how I like to work these, of course--if your way is easier for you and works, do that!

OpenStudy (hihi67):

you are very much a genius. that really helps!!!! you should become my math teacher!

OpenStudy (hihi67):

hah

OpenStudy (hihi67):

what about for e?

OpenStudy (hihi67):

e is 7/8m

OpenStudy (hihi67):

because: 1 1/2m is 3/2m and 3/4 m, normal. m/8 = 1/8 times m. 3/2 time m - 3/4 times m is 3/4 times m. 3/4 time m - 1/8 times m is 7/8

OpenStudy (agreene):

Its a little bit harder, but: I'd start by converting that mixed fraction in the front, 1 1/2 = 3/2 so 1 1/(2m) = 3/(2m) that gives us: \[\frac3{2m}-\frac3{4m}+\frac{m}8\] convert that 2m denominator to 4m by multiplying 2/2 through it: \[\frac6{4m}-\frac3{4m}+\frac{m}8\] 6-3 = 3 so: \[\frac3{4m}+\frac{m}8\] although you can simplify this a bit, id leave it there and call that my answer.

OpenStudy (hihi67):

okay..

OpenStudy (hihi67):

what i dont get, is f) how is -4a/5 + 1/10 a is -7/10a. its -9/10a!!!

OpenStudy (hihi67):

i solved it right, but for some reason i keep getting -9/10.

OpenStudy (hihi67):

which is sad, cause i should know how to solve that.

OpenStudy (agreene):

combining \[\frac{-4a}5+\frac1{10a}\] is a bit tricky, in fact in my first answer, i didn't even do it. the reason is the variable is in the numerator of one and the denominator of the other... if you would like to do it, remember you need to multiply through and get the same denominator... in this case: 10a is what we need. 5 is what we have, so we need to multiply through 2a to the fraction on the left: \[\frac{-4a}{5}*\frac{2a}{2a}+\frac1{10a}\] \[\frac{-8a^2}{10a}+\frac1{10a}\] now we can combine these two, since they have the same denominator: \[\frac{1-8a^2}{10a}\]

OpenStudy (hihi67):

oh. i just did 5 * 2 is 10. -4a * 2 is 8!

OpenStudy (agreene):

yeah this one is a bit more tricky :P

OpenStudy (hihi67):

okay. thank!

OpenStudy (hihi67):

you

OpenStudy (hihi67):

bye now!